# Cost-Free Neutrality for the River Method

### v1

@@ 1 Introduction and Related Work
A common interest in theoretical computer science, economics, and political science is the design and analysis of social choice functions-mechanisms that aggregate individual preferences to make a collective decision. Clearly, such decision-making is an integral part of democratic processes. Beyond political contexts, social choice functions have become increasingly relevant in artificial intelligence, particularly for aligning AI systems with diverse human preferences (furnkranz2003pairwise; kopf2024openassistant). As language models and other AI systems interact with feedback from a wide range of users, aggregating that feedback ‘socially acceptably’ becomes a key challenge. Recent work argues that social choice theory should guide AI alignment in addressing this diversity (conitzer_social_2024).
There is extensive research on properties of social choice functions, their trade-offs, and the often high complexity of computing them (see (brandt_handbook_2016) for an overview).
A core task of computational social choice is to design rules that satisfy reasonable fairness criteria while also being computationally tractable. In this work, we show how to efficiently compute the winners of the recently suggested River Method (doring_river_2025).
@@ 1.1 Margin-Based Social Choice Functions
The margin of $x$ over $y$ is the difference between the number of voters who prefer $x$ over $y$ and those who prefer $y$ over $x$. This information is represented as a margin graph – a complete, weighted, antisymmetric, directed graph over the set of alternatives, where each edge $(x,y)$ is weighted by the margin between $x$ and $y$ (brandt_handbook_2016). If there is no Condorcet winner, the margin graph contains cycles of pairwise victories, such as alternatives $x,y,z$ where $x$ wins against $y$, which wins against $z$, which wins against $x$.
@@
In every election there exists at least one immune alternative (holliday_split_2023). Among the numerous social choice functions selecting immune alternatives, we want to highlight five that are particularly notable for their distinct methods of resolving majority cycles:
  • Split Cycle (holliday_split_2023) aims to identify a consistent winner set by deleting the weakest link in every cycle in the margin graph. As Split Cycle selects all immune alternatives as winners, the following methods can be seen as refinements.
  • Ranked Pairs (tideman_independence_1987) incrementally locks in pairwise victories, prioritizing those with the largest margins that do not create cycles.
  • Stable Voting (holliday_stable_2023) recursively refines the winner set, ensuring stability under the deletion of alternatives.
  • Beat Path (schulze_new_2011; schulze_schulze_2025) selects winners based on the strongest paths through the defeat graph.
  • River (doring_river_2025) builds on Ranked Pairs by providing a simpler approach to refine the majority cycle elimination process.
For an overview of the definitions and properties of these methods, we refer to (holliday_split_2023; doring_river_2025).
@@
Unfortunately, such a tiebreaker usually compromises the property of neutrality, which requires that no alternative is favored a priori. See (brill_price_2012; freeman_general_2015; wang_practical_2019; zavist_complete_1989) for a discussion of possible tiebreaking rules and their implications. To reintroduce neutrality, one can compute the winning set using the Parallel-Universe Tiebreaking (PUT), which considers all possible tiebreaking orders and returns the union of winners. Unfortunately, deciding whether an alternative is a winner under PUT for Ranked Pairs is known to be $\mathtt{NP}$-complete (brill_price_2012).
@@ 2 Preliminaries
A Condorcet winner $x$ is an alternative which defeats all other alternatives in a pairwise comparison, i. e., $m(x,y)>0$ for all $y\in A\setminus\{x\}$. Such an alternative does not always exist. In the absence of a Condorcet winner, every alternative faces at least one majority defeat. Some alternatives can be justified as winners despite these defeats by the existence of equally strong majority paths leading from the alternative to each alternative defeating it. This notion is called immunity (holliday_split_2023): An alternative $x$ is immune (against majority complaints) if, for every alternative $y$ with $m(y,x)>0$, there is a majority path $P_{xy}$ with $\operatorname{strength}(P_{xy})\geq m(y,x)$.
@@
River (doring_river_2025) is a social choice function which eliminates majority cycles in the margin graph by greedily constructing a subtree of $\mathcal{M}$ based on a fixed tiebreaker. Formally, given a preference profile $\mathcal{P}$ and a tiebreaker $\tau\in\mathbb{L}$, let $\mathcal{M}=(A,E)$ denote the corresponding margin graph. River constructs the River diagram $\mathcal{M}^{\mathsf{RV}}(\mathcal{P},\tau)$, and selects a unique winner $\mathsf{RV}(\mathcal{P},\tau)$ as follows:
@@ 3.2 Computing the Fused-Universe Diagram
1
2
31ex
[…]
5
61exforeach $(x,y)\in E$ in order of decreasing margin do
7 $k\leftarrow m(x,y)$
8
@@ 7 $k\leftarrow m(x,y)$
9 if $y$ is fixedly dominated and $\exists(z,y)\in{\mathcal{M}^{\mathsf{FUN}}}\colon m(z,y)>k$ then reject $(x,y)$ and continue
10
@@
11 if there is a path from $y$ to $x$ of strength >k>k in $\mathcal{M}^{\mathsf{FUN}}$} then
12      compute the set of vertices $U$ that have a path to $x$ of strength >k>k in $\mathcal{M}^{\mathsf{FUN}}$}
13    if for all $u\in U$ there exists a path $P\in{\mathcal{M}^{\mathsf{FUN}}}$ from $y$ to $u$ of strength >k>k, such that for every CC edge $(c,d)\in P$, all paths from $d$ to $c$ in $\mathcal{M}^{\mathsf{FUN}}$} of strength $\geq m(c,d)$ include vertex $y$ then
14       reject $(x,y)$ and continue
15
16
17
@@
18 if $y$ is not dominated then $state(x,y)\leftarrow\textnormal{{Fix}}$
19 else if $y$ is cycle dominated and $\exists(z,y)\in{\mathcal{M}^{\mathsf{FUN}}}\colon m(z,y)>k$ then $state(x,y)\leftarrow\textsf{CBC}$
20 else if $y$ is cycle dominated and $\forall(z,y)\in{\mathcal{M}^{\mathsf{FUN}}}\colon m(z,y)=k$ then $state(x,y)\leftarrow\textsf{BC}$
21 else // $y$ is fixedly dominated by edges of margin $k$
22    $state(x,y)\leftarrow\textsf{BC}$
23      update incoming Fix edge of $y$ to BC
24
25 add $(x,y)$ to $\mathcal{M}^{\mathsf{FUN}}$}
26
@@ 22    $state(x,y)\leftarrow\textsf{BC}$
27  compute the set of edges $C\subseteq{\mathcal{M}^{\mathsf{FUN}}}$ that form a cycle with $(x,y)$
28 if $C\neq\emptyset$ then
29    $state(x,y)\leftarrow\textsf{CC}$
30    foreach $e\in C$ with $m(e)=k$ do $state(e)\leftarrow\textsf{CC}$
31
32
33return $\mathcal{M}^{\mathsf{FUN}}$}
@@ 4.2 Backward Direction: Cycle Edges and a Winning Certificate
The algorithm computing $T^{a}$ is a variant of Prim’s algorithm for computing a Minimum Spanning Tree (MST) (primShortestConnectionNetworks1957; jarnikJistemProblemuMinimalnim1930; dijkstraNoteTwoProblems1959), therefore named DirectedMaxPrim:
@@
An efficient implementation uses a priority queue such as Fibonacci Heaps (fredmanFibonacciHeapsTheir1987) to select the edges in step 2. For a $\mathsf{FUN}$ diagram with $|A|$ alternatives and $|E|$ edges, DirectedMaxPrim has a runtime in $O(|E|+|A|\log|A|)$ (cormenIntroductionAlgorithms2022).
@@ 5 Experiments
Figure 3: A scatterplot of all recorded running times of the $\mathsf{FUN}$ algorithm versus $\operatorname{\mathsf{RV\mbox{-}PUT}}$ and $\operatorname{\mathsf{RP\mbox{-}PUT}}$ as implemented in pref-voting by (mattei_preflib_2013) over synthetic election with different numbers of alternatives and voters on a logarithmic scale. Note that after 3 timeouts, no more running times where recorded for that $m$.
Figure 4: Average running time of the $\mathsf{FUN}$ algorithm versus Stable Voting, Beat Path, and Split Cycle as implemented in pref-voting by (mattei_preflib_2013) over single runs per 20 synthetic elections with different numbers of alternatives and voters on a logarithmic scale.
To assess the practical performance of the $\mathsf{FUN}$ algorithm, we conducted preliminary experiments comparing its running time to that of other voting methods. In particular, we evaluated River $\mathsf{FUN}$ against River with brute-force PUT ($\operatorname{\mathsf{RV\mbox{-}PUT}}$), Ranked Pairs with brute-force PUT ($\operatorname{\mathsf{RP\mbox{-}PUT}}$), as well as the polynomial-time methods Split Cycle, Stable Voting, and Beat Path, using implementations from the pref-voting Python library (mattei_preflib_2013).
@@ 5.1 Setup
We followed the recommendations by (szufa_drawing_2025) and generated synthetic elections using the normalized Mallows model with dispersion parameter $\phi=0.35$, which simulates elections that resemble real-world data. We focused only on elections without a Condorcet winner, since all considered methods are Condorcet consistent and first check for Condorcet winners, therefore taking the same amount of time to return the winner. We generated elections for all combinations of:
@@ 5.3 Results
$\mathsf{FUN}$ ’s performance is comparable to that of Split Cycle, Beat Path, and Stable Voting (Figure 4). The implementations of Split Cycle and Beat Path in (mattei_preflib_2013) both use the Floyd–Warshall algorithm, which explains their nearly identical performance. While $\mathsf{FUN}$ is slightly slower than the other methods for $m\leq 14$, its performance surpasses Stable Voting from $m\geq 15$ and surpasses/ is on par with Beat Path and Split Cycle from $m\geq 20$. Note that for $n\geq 50$, its running time closely matches that of Beat Path and Split Cycle.

### v2

@@ 1 Introduction and Related Work
A common interest in theoretical computer science, economics, and political science is the design and analysis of social choice functions-mechanisms that aggregate individual preferences to make a collective decision. Clearly, such decision-making is an integral part of democratic processes. Beyond political contexts, social choice functions have become increasingly relevant in artificial intelligence, particularly for aligning AI systems with diverse human preferences (Fürnkranz and Hüllermeier 2003; Köpf et al. 2024). As language models and other AI systems interact with feedback from a wide range of users, aggregating that feedback ‘socially acceptably’ becomes a key challenge. Recent work argues that social choice theory should guide AI alignment in addressing this diversity (Conitzer et al. 2024).
There is extensive research on properties of social choice functions, their trade-offs, and the often high complexity of computing them (see (Brandt et al. 2016) for an overview).
A core task of computational social choice is to design rules that satisfy reasonable fairness criteria while also being computationally tractable. In this work, we show how to efficiently compute the winners of the recently suggested River Method (Döring et al. 2025).
@@ 1.1 Margin-Based Social Choice Functions
The margin of $x$ over $y$ is the difference between the number of voters who prefer $x$ over $y$ and those who prefer $y$ over $x$. This information is represented as a margin graph – a complete, weighted, antisymmetric, directed graph over the set of alternatives, where each edge $(x,y)$ is weighted by the margin between $x$ and $y$ (Brandt et al. 2016). If there is no Condorcet winner, the margin graph contains cycles of pairwise victories, such as alternatives $x,y,z$ where $x$ wins against $y$, which wins against $z$, which wins against $x$.
@@
In every election there exists at least one immune alternative (Holliday and Pacuit 2023a). Among the numerous social choice functions selecting immune alternatives, we want to highlight five that are particularly notable for their distinct methods of resolving majority cycles:
  • Split Cycle (Holliday and Pacuit 2023a) aims to identify a consistent winner set by deleting the weakest link in every cycle in the margin graph. As Split Cycle selects all immune alternatives as winners, the following methods can be seen as refinements.
  • Ranked Pairs (Tideman 1987) incrementally locks in pairwise victories, prioritizing those with the largest margins that do not create cycles.
  • Stable Voting (Holliday and Pacuit 2023b) recursively refines the winner set, ensuring stability under the deletion of alternatives.
  • Beat Path (Schulze 2011, 2025) selects winners based on the strongest paths through the defeat graph.
  • River (Döring et al. 2025) builds on Ranked Pairs by providing a simpler approach to refine the majority cycle elimination process.
For an overview of the definitions and properties of these methods, we refer to (Holliday and Pacuit 2023a; Döring et al. 2025).
@@
Unfortunately, such a tiebreaker usually compromises the property of neutrality, which requires that no alternative is favored a priori. See (Brill and Fischer 2012; Freeman et al. 2015; Wang et al. 2019; Zavist and Tideman 1989) for a discussion of possible tiebreaking rules and their implications. To reintroduce neutrality, one can compute the winning set using the Parallel-Universe Tiebreaking (PUT), which considers all possible tiebreaking orders and returns the union of winners. Unfortunately, deciding whether an alternative is a winner under PUT for Ranked Pairs is known to be $\mathtt{NP}$-complete (Brill and Fischer 2012).
@@ 2 Preliminaries
A Condorcet winner $x$ is an alternative which defeats all other alternatives in a pairwise comparison, i. e., $m(x,y)>0$ for all $y\in A\setminus\{x\}$. Such an alternative does not always exist. In the absence of a Condorcet winner, every alternative faces at least one majority defeat. Some alternatives can be justified as winners despite these defeats by the existence of equally strong majority paths leading from the alternative to each alternative defeating it. This notion is called immunity (Holliday and Pacuit 2023a): An alternative $x$ is immune (against majority complaints) if, for every alternative $y$ with $m(y,x)>0$, there is a majority path $P_{xy}$ with $\operatorname{strength}(P_{xy})\geq m(y,x)$.
@@
River (Döring et al. 2025) is a social choice function which eliminates majority cycles in the margin graph by greedily constructing a subtree of $\mathcal{M}$ based on a fixed tiebreaker. Formally, given a preference profile $\mathcal{P}$ and a tiebreaker $\tau\in\mathbb{L}$, let $\mathcal{M}=(A,E)$ denote the corresponding margin graph. River constructs the River diagram $\mathcal{M}^{\mathsf{RV}}(\mathcal{P},\tau)$, and selects a unique winner $\mathsf{RV}(\mathcal{P},\tau)$ as follows:
@@ 3.2 Computing the Fused-Universe Diagram
1ex
${\mathcal{M}^{\mathsf{FUN}}}\leftarrow(A,\emptyset)$
1exforeach $(x,y)\in E$ in order of decreasing margin do
 $k\leftarrow m(x,y)$
@@
 if $y$ is fixedly dominated and $\exists(z,y)\in{\mathcal{M}^{\mathsf{FUN}}}\colon m(z,y)>k$ then reject $(x,y)$ and continue
@@
 if there is a path from $y$ to $x$ of strength >k>k in $\mathcal{M}^{\mathsf{FUN}}$} then
      compute the set of vertices $U$ that have a path to $x$ of strength >k>k in $\mathcal{M}^{\mathsf{FUN}}$}
    if for all $u\in U$ there exists a path $P\in{\mathcal{M}^{\mathsf{FUN}}}$ from $y$ to $u$ of strength >k>k, such that for every CC edge $(c,d)\in P$, all paths from $d$ to $c$ in $\mathcal{M}^{\mathsf{FUN}}$} of strength $\geq m(c,d)$ include vertex $y$ then
       reject $(x,y)$ and continue
      end if
   end if
@@
 if $y$ is not dominated then $state(x,y)\leftarrow\textnormal{{Fix}}$
 else if $y$ is cycle dominated and $\exists(z,y)\in{\mathcal{M}^{\mathsf{FUN}}}\colon m(z,y)>k$ then $state(x,y)\leftarrow\textsf{CBC}$
 else if $y$ is cycle dominated and $\forall(z,y)\in{\mathcal{M}^{\mathsf{FUN}}}\colon m(z,y)=k$ then $state(x,y)\leftarrow\textsf{BC}$
 else // $y$ is fixedly dominated by edges of margin $k$
    $state(x,y)\leftarrow\textsf{BC}$
      update incoming Fix edge of $y$ to BC
   end if
 add $(x,y)$ to $\mathcal{M}^{\mathsf{FUN}}$}
@@
  compute the set of edges $C\subseteq{\mathcal{M}^{\mathsf{FUN}}}$ that form a cycle with $(x,y)$
 if $C\neq\emptyset$ then
    $state(x,y)\leftarrow\textsf{CC}$
    foreach $e\in C$ with $m(e)=k$ do $state(e)\leftarrow\textsf{CC}$
   end if
end foreach
return $\mathcal{M}^{\mathsf{FUN}}$}
@@ 4.2 Backward Direction: Cycle Edges and a Winning Certificate
The algorithm computing $T^{a}$ is a variant of Prim’s algorithm for computing a Minimum Spanning Tree (MST) (Prim 1957; Jarník 1930; Dijkstra 1959), therefore named DirectedMaxPrim:
@@
An efficient implementation uses a priority queue such as Fibonacci Heaps (Fredman and Tarjan 1987) to select the edges in step 2. For a $\mathsf{FUN}$ diagram with $|A|$ alternatives and $|E|$ edges, DirectedMaxPrim has a runtime in $O(|E|+|A|\log|A|)$ (Cormen et al. 2022).
@@ 5 Experiments
Figure 3: A scatterplot of all recorded running times of the $\mathsf{FUN}$ algorithm versus $\operatorname{\mathsf{RV\mbox{-}PUT}}$ and $\operatorname{\mathsf{RP\mbox{-}PUT}}$ as implemented in pref-voting by (Mattei and Walsh 2013) over synthetic election with different numbers of alternatives and voters on a logarithmic scale. Note that after 3 timeouts, no more running times where recorded for that $m$.
Figure 4: Average running time of the $\mathsf{FUN}$ algorithm versus Stable Voting, Beat Path, and Split Cycle as implemented in pref-voting by (Mattei and Walsh 2013) over single runs per 20 synthetic elections with different numbers of alternatives and voters on a logarithmic scale.
To assess the practical performance of the $\mathsf{FUN}$ algorithm, we conducted preliminary experiments comparing its running time to that of other voting methods. In particular, we evaluated River $\mathsf{FUN}$ against River with brute-force PUT ($\operatorname{\mathsf{RV\mbox{-}PUT}}$), Ranked Pairs with brute-force PUT ($\operatorname{\mathsf{RP\mbox{-}PUT}}$), as well as the polynomial-time methods Split Cycle, Stable Voting, and Beat Path, using implementations from the pref_voting Python library (Holliday and Pacuit 2025).
@@ 5.1 Setup
We followed the recommendations by (Szufa et al. 2025) and generated synthetic elections using the normalized Mallows model with dispersion parameter $\text{norm-}\phi=0.7$, which simulates elections that resemble real-world data. We focused only on elections without a Condorcet winner, since all considered methods are Condorcet consistent and first check for Condorcet winners, therefore taking the same amount of time to return the winner. We generated elections for all combinations of:
@@ 5.3 Results
$\mathsf{FUN}$ ’s performance is comparable to that of Split Cycle, Beat Path, and Stable Voting (Figure 4). The implementations of Split Cycle and Beat Path in (Mattei and Walsh 2013) both use the Floyd–Warshall algorithm, which explains their nearly identical performance. While $\mathsf{FUN}$ is slightly slower than the other methods for $m\leq 16$, its performance surpasses Stable Voting from $m\geq 17$ and surpasses/ is on par with Beat Path and Split Cycle from $m\geq 20$. Note that for $n=10$, its running time closely matches that of Beat Path and Split Cycle.